OpenStudy (anonymous):

Compute and simplify the difference quotient f ( x + h ) − f ( x ) h for f ( x ) = 9 x − 11 .

OpenStudy (anonymous):

f ( x + h ) − f ( x )/ h - the entire thing is over h for f ( x ) = 9 x − 11 .

OpenStudy (anonymous):

Looking first at just the top (we'll deal with the h in the denominator at the end), f(x+h) - f(x) = [9(x + h) - 11] - [9x - 11] = 9x + 9h - 11 - 9x + 11 = 9h Now divide by h...

OpenStudy (anonymous):

thank you so much!